- Sat: Nakano, Enomoto, Guiguo, Yokozawa, Sugimoto
- Sun: Nakano, Enomoto, Guiguo, Yokozawa
Transition from morning work (*:responsible person)
- Current status as of 1/22: IR and Green reached X end PD.
1/22 Surface inspection for PR mirrors. Cleaning will be done after 1/31.
1/22 or 23 Surface inspection for SR mirrors.
- 1/24-25 Cleaning HR of SRs.
1/23 BS is aligned to Y-end PD by IR.
- 1/24-25 ITMY is aligned to REFL by IR.
- 1/24-25 ITMX is aligned to REFL by IR.
- 1/24-25 Beam position of IR at SR3 is confirmed how far from the center.
- 1/28-29 SR3 is aligned to SR2 center by IR.
- 1/30 Green beam is injected from SR2 AR surface to Y end PD.
- 1/31-2/1 Removing bellows at PR and SR area.
- 1/31-2/1 Cleaning HR of PR3 and AR of SR2
2/4-8 Alignment of SR2 -> SRM by IR, SRM -> OMMT by IR -> OMMT -> OSTM by IR, OSTM -> OMC by IR
2/4-8 Alignment of ETMY -> POS by Green
- 2/12-15 Closing door for central area, pumping down
- needs drift monitor for VIS
Report of work (*:responsible person)
- COM: IR beam reached at Y end
- We checked Y-end PDA
- Green beam for X arm was used for rough alignment of BS to center at ITMY
- Then IR beam hit the Y end.
- Type A:
- IP position of ITMY is somehow OK, but stepping motor of BF or is not working well.
- IP of ETMY is locked yesterday.
- TYPE B: GEO phone amp was replaced and it seems to be fine. We need a little more time to replace other GEO phone.
- MIR: HR of SRM cleaning is going on yesterday and today, HR of SR3 is the next.
- DGS: RFM network is available for both X and Y arm.
- CAM: k1cam0 was recovered.
- OMC: It was found that the beam is blocked by Type C OSEM. We need some modifications at the frame.
Work plan (*:responsible person)
- COM: Alignment of ITMY today.
- Type A: Releasing ETMY lock.